OpenStudy (anonymous):

find the range of y=2/x+2 A. all real numbers except y=2 B. y≥2 C. y≤2 D. y≥0

OpenStudy (blurbendy):

is that y = 2/ (x+2) or y = (2/x) + 2?

OpenStudy (anonymous):



OpenStudy (anonymous):

@blurbendy

OpenStudy (blurbendy):

A